# Direct view of gate-tunable miniband dispersion in graphene superlattices near the magic twist angle

### v1

@@
Figure 4: Bandwidth renormalization with doping in TBG and TDBG. a, LB and UB bandwidths obtained from the fitted TBG dispersions in Fig. 2(f). Dashed lines are linear fits that have been added as a guide to the eye. b, LB and UB bandwidths of TDBG extracted from the calculated spectral functions in Figs. 3(e) and Fig. 10 in the Supporting Information. The error bars for W in the TBG estimates are ±$\pm 19$ meV and ±$\pm 32$ meV for the LB and UB, respectively, which have been propagated from the dispersion fits in Fig. 9 in the Supporting Information. For TDBG the error bars on W are associated with how accurate we can determine the band edges, which leads to error bars smaller than the data points. The error bars on $\nu$ are (a) ±$\pm 0.52$ and (b) ±$\pm 0.64$ from the analysis in Figs. 6-7 in the Supporting Information, respectively.
@@
Our findings underline the vast design freedom in moiré systems by revealing a striking difference in the filling-dependent band dispersion of TBG and TDBG. The behaviour of TBG is non-trivial in the sense that the minibands experience a non-rigid shift upon increased filling, corresponding to a monotonous bandwidth change of the LB and UB in opposite directions upon increased filling (see Fig. 4(a)). However, this is easily captured by a simple tight-binding model with a variable hopping parameter $t$. Moreover, the results are consistent with the theoretically expected behaviour [8, 9, 11] and previous experiments on the single layer graphene [31]. By contrast, the filling-dependent band dispersion observed in TDBG is highly complex. The LB and UB bandwidth changes determined from the calculated spectral function are summarized in Fig. 4(b). Upon increasing the filling, all bands shift to lower energies but they also drastically change their dispersion with a non-monotonous bandwidth change of the LB. The $\nu$-dependence of the bandwidth ultimately influences the $U/W$ ratio in the LB and UB that leads to Mott insulating and superconducting phases at low temperatures. At $\nu=-2$ where the correlated insulator phase emerges in the LB we find bandwidths in the range of 10-20 meV for both TBG and TDBG. At $\nu=2$ where correlation effects emerge in the UB, we estimate somewhat larger bandwidths, on the order of 60 meV and 20 meV for TBG and TDBG, respectively. These values reflect the complex doping-dependent asymmetries between the flat band dispersions, which are important to consider for the theoretical understanding of doping-dependent many-body effects in the materials.
1 Methods
1.1 Fabrication of devices
@@
Stacking of 2D flakes was accomplished using the dry-transfer technique applying polymers [33, 34]. Graphene and hBN layers were exfoliated onto SiO2/Si substrates. Their size and thickness were identified by optical microscopy and AFM. Bilayer and monolayer graphene flakes were cut in two pieces by using a sharp tip with 2 $\mu$m diameter. One of the pieces was picked up using a polycarbonate (PC)/polydimethylsiloxane (PDMS) stamp mounted on a glass slide at 130 C. The remaining piece was rotated by an angle $\theta$ [16], which was set $0.1^{\circ}$ higher than the target angle and stacked with the mono- or bilayer graphene flake on the stamp in order to form twisted bilayer graphene or twisted double bilayer graphene. Then, a hBN flake with thickness around $30$ nm was picked up at 120 C, which was followed by picking up a graphite flake at 130 C to be used as a back gate. Finally, the stack was released onto SiO2/Si with Au/Cr-patterned electrodes based on $50$ nm gold and $5$ nm chromium films at 180 C. Twisted graphene layers and back gates were contacted by different electrodes. Heterostructures were then annealed in a H2/Ar environment at 250 C for several hours prior to measurement.
1.2 C-AFM Measurements
@@
1.3 microARPES measurements
The microARPES measurements were performed at the SGM4 beamline of the ASTRID2 light source at Aarhus University, Denmark [35]. The devices were wirebonded to CSB00815 chip carriers and annealed in the microARPES end-station at a temperature of 180 °$\degree$°C for 1 hour before exposure to the beam. Throughout the data collection, the samples were kept at room temperature in a base pressure better than $2\cdot 10^{-10}$ mbar.
@@
1.4 Spectral function calculations for TDBG
Fully relaxed crystal structure of TDBG were obtained on the molecular dynamics simulation using LAMMPS [36] software with REBO2 [37] and exact exchange random phase approximation (EXX-RPA)-fitted DRIP potentials [38]. The relaxation was performed for a $1.35^{\circ}$ twisted commensurate cell made with two AB stacked bilayer graphene with a lattice constant of $a=2.46\text{\AA}$. We performed calculations for 14,408 carbon $\pi$-orbitals within the relaxed structure of moiré lattice length of $104.39$.
@@ 1.4 Spectral function calculations for TDBG
Here, $\Delta\rho$ denotes the disparity between the density matrices of $H$ and the non-corrected Hamiltonian $H^{0}=\sum_{ij\sigma}t_{ij}c^{\dagger}_{i\sigma}c_{j\sigma}$. The operator $c_{i\sigma}$($c^{\dagger}_{i\sigma}$) is electron annihilation(creation) operators, where $\sigma$ denotes the spin index, and $\tilde{\sigma}$ signifies the spin in the opposite direction. The hopping parameters $t_{ij}$ were determined using the scaled hybrid exponential (SHE) model, which incorporates the effects of structure relaxation [38]. We assumed the onsite Hubbard parameter $U=5$ eV. For intersite interactions $V_{ij}$, values of $V_{1}=2$ eV for first nearest neighbors and $V_{2}=1$ eV for second nearest neighbors were assumed in intralayer relations, while for interlayer relations, the Coulomb equation accounting for bonding length was employed, as $V_{ij}=1/(\epsilon_{r}\sqrt{(a/2\sqrt{3})^{2}+r_{ij}^{2}})$ with the relative permittivity $\epsilon_{r}=6$ [40]. The exchange interaction was disregarded to simplify the calculations, hence the calculations were carried out at the self-consistent Hartree level. Each gate voltage was considered to be applied to tDBG on 30nm hBN, with saw-like external potential and doping corresponding to the same dielectric constant. The Fermi level of the calculations has been matched to the experiment by comparing the integrated spectral weight with the Lorentzian peaks obtained from the EDC analysis in Fig. 3(c), as shown in Fig. 11 in the Supporting Information. The self-consistent calculations were performed with $5\times 5$ grid k-points.

### v2

@@ Discussion
Figure 4: Bandwidth renormalization with doping in TBG and TDBG. a, LB and UB bandwidths obtained from the fitted TBG dispersions in Fig. 2(f). Dashed lines are linear fits that have been added as a guide to the eye. b, LB and UB bandwidths of TDBG extracted from the calculated spectral functions in Figs. 3(e) and Fig. 10. The error bars for the LB bandwidth in TBG have been propagated from the dispersion fits in Fig. 9, while the error bars on the UB bandwidth are propagated from the LB bandwidth and $n$. For TDBG the error bars on W are associated with how accurate we can determine the band edges, which leads to error bars smaller than the data points. The error bars on $n$ are derived from the calibrations shown in Figs. 6-7.
@@
Our findings underline the vast design freedom in moiré systems by revealing a striking difference in the doping-dependent band dispersion of TBG and TDBG. The behaviour of TBG is non-trivial in the sense that the minibands experience a non-rigid shift upon increased filling, corresponding to a monotonous bandwidth change of the LB and UB in opposite directions (see Fig. 4(a)). However, this is easily captured by a simple tight-binding model with a variable hopping parameter $t$. Moreover, the results are consistent with the theoretically expected behaviour [15, 16, 18] and previous experiments on single layer graphene [29]. By contrast, the filling-dependent band dispersion observed in TDBG is highly complex. The LB and UB bandwidth changes determined from the calculated spectral function are summarized in Fig. 4(b). Upon increasing the filling, all bands shift to lower energies but they also drastically change their dispersion with a non-monotonous bandwidth change of the LB. The $n$-dependence of the bandwidth ultimately influences the $U/W$ ratio in the LB and UB that leads to Mott insulating and superconducting phases at low temperatures. At doping levels where the correlated insulator phase emerges in the LB we find bandwidths in the range of 10-20 meV for both TBG and TDBG. In the doping regime where correlation effects emerge in the UB, we estimate somewhat larger bandwidths, on the order of 60 meV and 20 meV for TBG and TDBG, respectively. These values reflect the complex doping-dependent asymmetries between the flat band dispersions that underpin the rich temperature- and doping-dependent electronic phase diagrams in these materials [31].
Methods
@@
Fabrication of devices. Stacking of 2D flakes was accomplished using the dry-transfer technique applying polymers [32, 33]. Graphene and hBN layers were exfoliated onto SiO2/Si substrates. Their size and thickness were identified by optical microscopy and AFM. Bilayer and monolayer graphene flakes were cut in two pieces by using a sharp tip with 2 $\mu$m diameter. One of the pieces was picked up using a polycarbonate (PC)/polydimethylsiloxane (PDMS) stamp mounted on a glass slide at 130 C. The remaining piece was rotated by an angle $\theta$ [4], which was set $0.1^{\circ}$ higher than the target angle and stacked with the mono- or bilayer graphene flake on the stamp in order to form twisted bilayer graphene or twisted double bilayer graphene. Then, a hBN flake with thickness around $30$ nm was picked up at 120 C, which was followed by picking up a graphite flake at 130 C to be used as a back gate. Finally, the stack was released onto SiO2/Si with Au/Cr-patterned electrodes based on $50$ nm gold and $5$ nm chromium films at 180 C. Twisted graphene layers and back gates were contacted by different electrodes. Heterostructures were then annealed in a H2/Ar environment at 250 C for several hours prior to measurement.
@@
microARPES measurements. The microARPES measurements were performed at the SGM4 beamline of the ASTRID2 light source at Aarhus University, Denmark [34]. The devices were wirebonded to CSB00815 chip carriers and annealed in the microARPES end-station at a temperature of 180 °$\degree$°C for 1 hour before exposure to the beam. Throughout the data collection, the samples were kept at room temperature in a base pressure better than $2\cdot 10^{-10}$ mbar.
@@
Spectral function calculations for TDBG. Fully relaxed crystal structure of TDBG were obtained on the molecular dynamics simulation using LAMMPS [35] software with REBO2 [36] and exact exchange random phase approximation (EXX-RPA)-fitted DRIP potentials [37]. The relaxation was performed for a $1.35^{\circ}$ twisted commensurate cell made with two AB stacked bilayer graphene with a lattice constant of $a=2.46\text{\AA}$. We performed calculations for 14,408 carbon $\pi$-orbitals within the relaxed structure of moiré lattice length of $104.39$.
@@
Here, $\Delta\rho$ denotes the disparity between the density matrices of $H$ and the non-corrected Hamiltonian $H^{0}=\sum_{ij\sigma}t_{ij}c^{\dagger}_{i\sigma}c_{j\sigma}$. The operator $c_{i\sigma}$($c^{\dagger}_{i\sigma}$) is electron annihilation(creation) operators, where $\sigma$ denotes the spin index, and $\tilde{\sigma}$ signifies the spin in the opposite direction. The hopping parameters $t_{ij}$ were determined using the scaled hybrid exponential (SHE) model, which incorporates the effects of structure relaxation [37]. We assumed the onsite Hubbard parameter $U=5$ eV. For intersite interactions $V_{ij}$, values of $V_{1}=2$ eV for first nearest neighbors and $V_{2}=1$ eV for second nearest neighbors were assumed in intralayer relations, while for interlayer relations, the Coulomb equation accounting for bonding length was employed, as $V_{ij}=1/(\epsilon_{r}\sqrt{(a/2\sqrt{3})^{2}+r_{ij}^{2}})$ with the relative permittivity $\epsilon_{r}=6$ [39]. The exchange interaction was disregarded to simplify the calculations, hence the calculations were carried out at the self-consistent Hartree level. Each gate voltage was considered to be applied to tDBG on 30nm hBN, with saw-like external potential and doping corresponding to the same dielectric constant. The Fermi level of the calculations has been matched to the experiment by comparing the integrated spectral weight with the Lorentzian peaks obtained from the EDC analysis in Fig. 3(c), as shown in Fig. 11. The self-consistent calculations were performed with $5\times 5$ grid k-points.